OpenStudy (anonymous):

what are foci of the the graph 4x^2 - 9y^2 = 36

OpenStudy (jdoe0001):

\(\bf 4x^2 - 9y^2 = 36 \implies \cfrac{4x^2}{36}-\cfrac{ 9y^2}{36} = 1\)

OpenStudy (jdoe0001):

\(\bf \cfrac{\cancel{4}x^2}{\cancel{36}}-\cfrac{ \cancel{9}y^2}{\cancel{36}} = 1 \implies \cfrac{x^2}{9}-\cfrac{ y^2}{4} = 1\)

OpenStudy (jdoe0001):

that means that your a^2 = 9, b^2 = 4 notice the center, is at the origin also keep in mind that the distance from the center is "c" and \(\bf c^2 = a^2+b^2\)

OpenStudy (jdoe0001):

also notice that the fraction with the POSITIVE sign, is the one with the "x" variable that means is a horizontal traverse axis, so it's opening to the left and right so find "c" distance, and move left and right THAT MUCH from (0, 0)

OpenStudy (anonymous):

so the answer would be ?

OpenStudy (jdoe0001):

so, what did you get for "c"?

OpenStudy (anonymous):

(0,3.6)

OpenStudy (jdoe0001):

hmm I got something else :/

OpenStudy (anonymous):

wat you get?

OpenStudy (jdoe0001):

actaully nope, you're correct

OpenStudy (jdoe0001):

I was getting dual squared heheh

OpenStudy (jdoe0001):

was squaring the 9, but 9 is already squared :(